Presented below is information related to Shawshank Real Estate Agency. Oct. 1 Pete Shawshank begins business as a real estate agent with a cash investment of $25,000 in exchange for common stock. 2 Hires an administrative assistant. 3 Purchases office furniture for $2,900, by paying $700 cash with the balance on account. 6 Sells a house and lot for N. Kidman, earning a fee of $3,600, with $600 collected in cash and the balance billed to N. Kidman. 27 Pays $900 on the balance related to the transaction of October 3. 30 Pays the administrative assistant $2,300 in salary for October. 31 Collects $1,200 of the balance owed by N. Kidman. Instructions: (a) Journalize the transactions. (You may omit explanations.) (b) What balance would Shawshank Real Estate Agency report for Accounts Payable in its October 31 financial statements? In which category of which financial statements would it be found? (c) What balance would Shawshank Real Estate Agency report for Accounts Receivable in its October 31 financial statements? In which category of which financial statements would it be found?
Answer #1

(a) The journal entries are provided below:

S. No Date Particulars Debit Credit
1 Oct 1 Cash 25,000.00
Common stock 25,000.00
2 Oct 3 Office furniture 2,900.00
Cash 700.00
Accounts payable 2,200.00
3 Oct 6 Cash 600.00
Accounts receivable 3,000.00
Revenue 3,600.00
4 Oct 27 Accounts payable 900.00
Cash 900.00
5 Oct 30 Salary 2,300.00
Cash 2,300.00
6 Oct 31 Cash 1,200.00
Accounts receivable 1,200.00

(b): Accounts payable balance on 31st October = 2200-900 = 1,300. This will be found as a current liability in the liabilities section of the balance sheet.

(c): Accounts receivable balance on 31st October = 3000-1200 = 1,800. This will be found as a current asset in the assets section of the balance sheet.